OpenStudy (anonymous):

which of the following equations represent the greatest value of x? x^2=17 x^2=26 x^3=64 x^3=27 PLESE HELP IM GOING TO GET KICKED FROM MY CLASS IF ID DONT GET THIS DONE TODAY

OpenStudy (anonymous):

please*

OpenStudy (anonymous):

not really clear what this means perhaps you are supposed to solve each for x and see what it the biggest solution?

OpenStudy (sleepyhead314):

we know that square root of 16 = 4 so x^2 = 17 would mean that x is close to 4

OpenStudy (sleepyhead314):

same idea with the second one 26 is close to 25 sqrt of 25 = 5 so x^2 = 26 means that x is close to 5

OpenStudy (anonymous):

actually if \(x^2=26\) then \(x=\pm\sqrt{26}\) but i suppose you are supposed to pick that one anyway

OpenStudy (sleepyhead314):

to compare them I was attempting to make the numbers look friendlier, but yeah that's the more accurate form

OpenStudy (anonymous):

Let's re-arrange and solve \[\large x^{2}=17 \rightarrow x=\sqrt{17}\]\[\large x^{2}=26 \rightarrow x=\sqrt{26}\]\[\large x^{3}=64 \rightarrow x=\sqrt[3]{64}\]\[\large x^{3}=27 \rightarrow x=\sqrt[3]{27}\]